Find the x- and y- intercepts
5x+y=15

Found 2 solutions by palanisamy, stanbon:
Answer by palanisamy(496)

The equation of the line is 5x+y = 15 ..(1)
To find the x-intercept, put y=0 in (1)
We get 5x=15
x=15/5
x=3
So, the x-intercept is (3,0)
To find the y-intercept, put x=0 in (1)
We get y=15
So, the y-intercept is (0,15)

Answer by stanbon(75887)

Find the x- and y- intercepts
5x+y=15
--------------
x-intercept?
Let y = 0 and solve for "x":
5x = 15
x = 3
--------------
y-intercept?
Let x = 0 and solve for "y":
y = 15
====================
Cheers,
Stan H.